OpenStudy (anonymous):

Find radius and center (x+4)^2+(y-6)^2=64

OpenStudy (anonymous):

the center of the circle is the point (-4,6) and the radius is sqrt(64)=8 because the center-radius form of the circle equation is in the format (x – h)^2 + (y – k)^2 = r^2 in your case h=-4 and k=6 and r^2=64

OpenStudy (anonymous):

right (y)